OpenStudy (anonymous):

Select the number sentence that illustrates the associative property of addition

OpenStudy (anonymous):

Select the number sentence that illustrates the associative property of addition.Select one of the options below as your answer:A. (5 × 4) + 10 = 10 × (5 + 4) B. (5 + 4) + 10 = 5 + (4 + 10) C. (5 + 4) × 10 = 5 + (10 × 4) D. (5 × 4) + 10 = 5 × (10 + 5)